OpenStudy (anonymous):

Given the function g(x) = 8x - 2, find g(-2).

OpenStudy (anonymous):

Substitute in the function the input term. If g(x) = 8*x-2, than g(y) = 8*y-2, g(1) = 8*1-2 and so on...

OpenStudy (anonymous):

put value of x and solve

OpenStudy (anonymous):

\[g(-2)=8(-2)-2\]solve

OpenStudy (anonymous):

so it would be -18 rite?